OpenStudy (lizilizi28):

SOLVE AND CHECK 27X^2-20X=0

OpenStudy (anonymous):

x(27x-20)=0 so x=0 or 27x=20 x=20/27

OpenStudy (anonymous):

Factor out x: \[x (27x - 20) = 0\] This is true if \[x=0\] or \[27x-20=0\]Solve these for x and check.